OpenStudy (anonymous):

Will fan and medal The distance traveled by a falling object is given by the formula d = 0.5gt2 where d = distance, g = the acceleration due to gravity, and t = time. Solve this equation for g, and use your formula to determine the acceleration due to gravity if a baseball takes 10 seconds to hit the ground after being dropped from a height of 490 meters. Show all steps in your work.

OpenStudy (anonymous):

what part are you having trouble with?

OpenStudy (anonymous):

Work: d = 0.5gt^2 d/0.5t^2 = g g = d/0.5t^2 t = 10 sec d = 490 ft Answer: g = 9.8 ft/sec² Is it right?

OpenStudy (anonymous):

Sorry was gone forgot I posted this xD I will let you know if it is right when i turn it in